# Unlocking Personalized Knowledge in Federated Large Language Model: The Power of Mixture of Experts

### v1

@@ Abstract
The Mixture of Experts (MoE) architecture has emerged as a prominent strategy for scaling large language models (LLMs), effectively leveraging sparse activation and facilitating task-specific personalization. However, current federated learning (FL) approaches are primarily designed for dense models, making them unable to directly exploit the sparsity inherent in MoE architectures. Treating MoE models as dense networks in federated scenarios results in excessive communication overhead and computational costs, undermining the potential for personalized knowledge sharing. To address these challenges, we propose FLEx (Federated LLMs with Personalized Experts), a novel federated learning framework explicitly tailored for MoE-based LLMs. FLEx efficiently personalizes by pruning the global MoE model to keep only one expert per client, and employs an adaptive gating mechanism to reintegrate these personalized experts into the pre-trained MoE layers, ensuring the original backbone architecture remains unchanged. These personalized experts are trained with local data and stored locally on each client, while the shared modules are aggregated globally. Extensive evaluations on diverse instruction-based datasets under non-IID conditions consistently demonstrate that FLEx outperforms existing federated baselines. Our code is available at https://anonymous.4open.science/r/FLEx-8F12.
@@ 1 Introduction
Large language models (LLMs) have seen a remarkable increase in both scale and capability in recent years, largely due to the rise of the Mixture-of-Experts (MoE) paradigm jacobsAdaptiveMixturesLocal1991 ; fedusSwitchTransformersScaling2022 ; jiangMixtralExperts2024 . By activating only a small subset of expert sub-networks per token, MoE architectures separate the number of parameters from the computational cost, enabling state-of-the-art performance at significantly lower inference overhead compared to dense Transformers of similar size. Recent advances deepseek-aiDeepSeekV3TechnicalReport2025 ; qwen2.5 highlight that MoEs are not just an efficient compression strategy, but a powerful means of task-specific personalization, where experts naturally specialize in different linguistic styles, domains, or user preferences.
Meanwhile, federated learning (FL) has become a compelling framework for training and adapting LLMs on privacy-sensitive data that remains on users’ devices zhangBuildingFederatedgptFederated2024 ; yeOpenFedLLMTrainingLarge2024 ; longDualpersonalizingAdapterFederated2024 ; wangFLoRAFederatedFineTuning2024 . However, nearly all existing FL algorithms for LLMs are designed for dense model architectures. Applying them directly to MoE models requires each client to transmit all experts, even those unrelated to its local data, resulting in high communication costs and diminished personalization. Early attempts to bridge MoE and FL zouFedMoEEfficientFederated2025 ; meiFedMoEPersonalizedFederated2024 ; luoMixtureExpertsMade2025 were developed for traditional models rather than trillion token pretrained LLMs. While they prove that federated MoE training is feasible, they still require each client to manage and transmit multiple experts, including those that are not activated by local data, resulting in unnecessary communication overhead, which makes these methods economically impractical and leaves a significant efficiency gap in federated training of large-scale pretrained LLMs.
Deploying MoE-based LLMs in a federated setting introduces two main challenges. First, the communication and computation costs that arise from the sparse structure of MoE models are high, especially under limited bandwidth and constrained local device resources. Second, existing methods primarily focus on personalized experts trained from local data while neglecting global information, failing to achieve an effective balance between personalization and global knowledge within MoE-based LLMs.
To address these challenges, we propose FLEx (Federated LLMs with Personalized Experts), a simple framework that combines the sparsity of Mixture-of-Experts models with the benefits of federated personalization. In FLEx, each client selects and fine-tunes only one expert from the global MoE model, tailored to its local data. During fine-tuning, a lightweight gating module integrates the personalized expert’s output into the original MoE layer, allowing the model to leverage expert specialization without altering the backbone architecture. In FLEx, clients locally update their chosen expert and specific parameters in the shared module. Only the shared module updates are then communicated for aggregation into the global model, ensuring shared improvements, while the expert’s modifications remain local to the client for personalization. Our main contributions are as follows:
  • We propose FLEx, a novel federated learning framework explicitly tailored for MoE-based large language models. FLEx addresses the lack of federated learning approaches specifically designed for sparse MoE architectures, effectively balancing global knowledge sharing and local personalization, thereby significantly reducing communication and computational costs compared to traditional FL methods.
  • FLEx efficiently personalizes by pruning the MoE model to retain a single expert per client and using an adaptive gating mechanism to reintegrate personalized experts into the pre-trained MoE layers without altering the original backbone architecture. These personalized experts are trained using local data and stored locally on each client, while the shared modules are aggregated globally. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
  • Extensive evaluations conducted on multiple challenging instruction-based datasets, including Databricks-dolly-15k, Alpaca-GPT4, Finance-Alpaca, and MedAlpaca, under non-IID scenarios demonstrate that FLEx consistently outperforms existing federated learning baselines across diverse tasks.
@@ 2.1 Federated Learning on Large Language Models
Federated Learning (FL) has gained significant attention as a promising method for training Large Language Models (LLMs), addressing challenges such as decentralized data management and privacy concerns zhangBuildingFederatedgptFederated2024 ; yeOpenFedLLMTrainingLarge2024 ; wangFLoRAFederatedFineTuning2024 . Recent surveys yaoFederatedLargeLanguage2024 ; yeOpenFedLLMTrainingLarge2024 have identified important intersections between FL and LLMs, particularly in areas such as efficient foundation model training, federated fine-tuning techniques, and the collaborative potential of FL in advancing LLM development. Notable contributions in this domain include FedIT zhangBuildingFederatedgptFederated2024 , which demonstrated the utility of FL in instruction tuning tasks for LLMs. Further innovations involve personalized LLM training methods like DualLoRA longDualpersonalizingAdapterFederated2024 and FDLoRA qiFDLoRAPersonalizedFederated2024 , which combine federated learning with personalized adapter methods. Other approaches such as FRLoRA yanFederatedResidualLowRank , FLoRA wangFLoRAFederatedFineTuning2024 , and FlexLoRA baiFederatedFinetuningLarge introduce new aggregation strategies designed to optimize low-rank fine-tuning and allow more flexible model updates in federated settings. Additionally, selective aggregation methods, like those presented in guoSelectiveAggregationLowRank2025 , further improve federated LLM training by refining the aggregation process.
@@ 2.2 Mixture of Experts
Mixture-of-Experts (MoE) jacobsAdaptiveMixturesLocal1991 ; fedusSwitchTransformersScaling2022 models activate only a subset of expert layers during inference or training, which significantly reduces computational costs while maintaining high performance. In recent, MoE architectures have gained attention for their scalability and efficiency in large language models jiangMixtralExperts2024 ; deepseek-aiDeepSeekV3TechnicalReport2025 ; qwen2.5 .
However, existing federated MoE approaches zhanFedMoEDAFederatedMixture2024 ; luoMixtureExpertsMade2025 are primarily designed for dense model architectures. Applying these methods directly to MoE models can result in substantial computational and communication overhead. To address this limitation, we propose FLEx, a novel framework specifically designed for federated training of MoE-based LLMs.
@@ 2.3 Expert Pruning for LLMs
Model pruning is a widely used technique to improve communication and computation efficiency in deep learning. In the context of MoE-based LLMs, recent pruning methods, such as Wanda sunSimpleEffectivePruning2024 and SparseGPT frantarSparsegptMassiveLanguage2023 , offer effective strategies to reduce the number of model parameters. Additionally, recent studies luNotAllExperts2024 ; liuEfficientExpertPruning2024 have demonstrated that many experts are either unimportant or redundant during inference on specific benchmarks. However, in federated learning with MoE-based LLMs, directly pruning certain experts may result in the loss of essential foundational knowledge. To address this, we propose pruning personalized side experts on local data. Furthermore, we integrate the pruned experts back into the original pretrained model to achieve an optimal balance between local adaptation and global generalization.
@@ 3 Preliminary
We consider a federated learning (FL) setup with $K$ clients, each holding a private dataset $\mathcal{D}_{k}$. In this setup, each client independently trains a model on its local data, while the central server periodically aggregates the updates to improve the global model. A widely adopted approach, the Federated Averaging (FedAvg) algorithm, aggregates the global parameters by computing a weighted average of the local parameters:
$W^{t+1}=\sum_{i=1}^{N}\frac{n_{i}}{n}W_{i}^{t+1},$ (1)
where $n=\sum_{i=1}^{N}n_{i}$ is the total number of training samples across all clients, and $W^{t+1}_{i}$ and $W^{t+1}$ are weight of $i$-th client and server in $t+1$-th round. This approach enables the global model to leverage distributed data while maintaining privacy.
@@
The Mixture of Experts (MoE) architecture introduces a sparse activation mechanism, where only a subset of experts is activated for each input. This design enables large-scale models to scale efficiently while reducing computational costs. In modern LLMs, the backbone consists of transformer layers, each comprising a self-attention layer and a feedforward network (FFN) layer:
[…]
$h_{t}^{l}$ $=\text{FFN}\left(u_{t}^{l}\right)+u_{t}^{l},$ (3)
where $u_{1:T}^{l}$ is hidden states of all tokens after the $l$-th attention module and $h_{t}^{l}$ is the output hidden state of the $t$-th token after the $l$-th Transformer block. In MoE-based models, the FFN layer is replaced by a MoE layer, which consists of multiple experts. This substitution yields:
$h_{t}^{l}=\sum_{i=1}^{N}\left(g_{i,t}\text{FFN}_{i}\left(u_{t}^{l}\right)\right)+u_{t}^{l},$ (4)
where $N$ denotes the total number of experts, and $g_{i,t}$ is the gating function that selects which experts are activated for each token $t$. The gating function $g_{i,t}$ is defined as:
$g_{i,t}=\begin{cases}s_{i,t},&\text{if }s_{i,t}\in\text{Topk}\left(\{s_{j,t}|1\leq j\leq N\},K\right)\\ 0,&\text{otherwise},\end{cases}$ (5)
where $s_{i,t}$ is the relevance score for expert $i$ for token $t$, computed using the softmax function:
$s_{i,t}=\text{Softmax}_{i}\left(\text{router}(u_{t})\right).$ (6)
Here, the router function $\text{router}(u_{t})$ computes the raw scores for each expert based on the token’s representation $u_{t}$, and the softmax function $\text{Softmax}_{i}$ normalizes these scores to determine the relevance of each expert. The gating mechanism activates only the top $K$ experts for each token $t$, which significantly reduces computational costs. However, if we treat MoE-based LLMs as dense models and aggregate the entire model, the communication cost would grow significantly, making it impractical. To address this, we propose FLEx and unlock the power for experts in LLMs.
@@ 4 Unlock the Power of Experts: FLEx
Federated learning (FL) enables distributed model training across multiple clients, each holding a private dataset. Mixture of Experts (MoE) models employ sparse activation, activating only a small subset of experts for each input. This sparsity makes it impractical to aggregate the entire model across all clients due to significant communication overhead. In this paper, we propose FLEx, a federated learning approach that optimizes MoE models by pruning the MoE layer to retain a personalized expert for each client. FLEx reduces communication costs, maintains high performance, and preserves knowledge from the original pre-trained MoE model. By integrating personalized experts into each client’s model, FLEx strikes a balance between global model benefits and local data needs. The framework of FLEx is shown in Figure 1.
@@ 4.1 Selective Aggregation in FLEx
In MoE-based LLMs, communication efficiency is crucial. FedAvg, the standard federated training algorithm, faces challenges when aggregating both dense attention layers and sparse MoE layers. While MoE layers activate only a subset of experts for each input, the attention layers in LLMs are dense and require substantial resources for aggregation in a federated setup. To address this, we propose a hybrid approach. The attention layers remain shared and aggregated globally, while the MoE layers are pruned locally. Instead of aggregating all MoE parameters, we fix the MoE layer’s weights and prune it to retain a personalized expert for each client.
4.2 Personalized Experts via MoE Pruning
To further enhance personalization and computational efficiency in federated learning, inspired by luNotAllExperts2024 , we introduce a pruning strategy based on MoE layers. This approach allows each client to select the most suitable subset of experts for their local data, maximizing personalization while avoiding the redundancy and resource consumption that may arise during global aggregation. Next, we will detail the specific steps and optimization strategies involved in this process.
Each client $i$ computes the output of all experts for a given input $x\in\mathcal{D}_{i}$. Let $F^{(l)}(x)$ denote the output of the MoE layer, and let $S\subset\{1,\dots,K\}$ represent the subset of experts retained after pruning. The pruned output at layer $l$ is given by $F^{\prime(l)}(x,S)$.
For personalization, each client selects the subset of experts that best approximates the original MoE output for their local dataset at each layer by minimizing the reconstruction loss, as shown in Step 1 of Figure 1. This optimization problem for a given layer $l$ is formulated as:
$S_{i}^{(l)}=\arg\min_{S\subset\{1,\dots,K\},|S|=n}\left(\frac{1}{|\mathcal{D}_{i}|}\sum_{x\in\mathcal{D}_{i}}\|F_{S}^{(l)}(x)-F^{(l)}(x)\|_{F}\right).$
To simplify the problem, we reduce the search space by setting $n=1$, choosing only a single expert per client. This reduces the computational complexity while ensuring effective personalization. Through this optimization problem, we greedyly select the personalized side experts at each layer.
4.3 Personalized Layer Integration
Once the personalized expert is selected, it is integrated into the global model by combining it with the global MoE layer, as shown in Step 2 of Figure 1. During the federated aggregation process, the LoRA adapter for the attention layer is aggregated globally, while the LoRA adapter for the personalized expert remains on the client side. This approach ensures that each client benefits from both the global model’s knowledge and the personalized expert’s adaptation to local data. Then the forward equation of MoE layer (4)
$h_{t}^{l}$ $=\sum_{i=1}^{N}\left(g_{i,t}\text{FFN}_{i}\left(u_{t}^{l}\right)\right)+g_{e}^{t}\text{FFN}_{e}(u_{t}^{l})+u_{t}^{l},$ (7)
$g_{e}^{t}$ $=\text{sigmoid}(\text{router}_{e}(u_{t}^{l})),$ (8)
where $\text{router}_{e}$ and $\text{FFN}_{e}$ denote the personalized router and personalized side expert generated by pruned experts, respectively. Here, we keep the original FFN layers fixed to ensure that our method preserves as much knowledge as possible for trillion-token pretrained MoE-based LLMs.
Remark.
Our method retains the pre-trained weights in the MoE layer while adding a personalized side expert. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.

### v2

# FLEx: Personalized Federated Learning for Mixture-of-Experts LLMs via Expert Grafting

@@ Abstract
Federated instruction tuning of large language models (LLMs) is challenged by significant data heterogeneity across clients, demanding robust personalization. The Mixture of Experts (MoE) architecture, where experts can specialize in distinct data patterns, presents a natural architectural solution to this challenge. The inherent sparsity of the MoE architecture, achieved by selectively activating experts, poses a significant challenge to its integration with federated learning (FL). Conventional FL frameworks, designed for dense models, naively aggregate all expert parameters irrespective of their local activation patterns. This naive approach not only undermines MoE’s dynamic sparsity but also risks corrupting the world knowledge within pretrained experts. To address this, we propose FLEx (Federated LLMs with Personalized Experts), a novel framework that leverages pretrained MoE-based LLMs for efficient personalization. By aggregating only the shared non-expert parameters, FLEx significantly reduces communication overhead and preserves the world knowledge stored within the frozen pretrained experts. For personalization, we introduce a novel expert grafting mechanism that leverages dynamic sparsity to construct a client-specific expert from selected components of pretrained experts, tailored to local data. This grafted expert is then fine-tuned locally alongside the gating mechanism. This joint training enables the model to learn when to leverage the shared knowledge from frozen experts and when to employ the personalized one. Evaluations on diverse, non-IID instruction tuning datasets show that FLEx consistently outperforms federated baselines on average, while demonstrating strong knowledge preservation on the knowledge-driven benchmark MMLU. Our code is available at https://anonymous.4open.science/r/FLEx-8F12.
@@ 1 Introduction
The widespread adoption of Large Language Models (LLMs) makes fine-tuning on decentralized user data essential for personalization. Federated Learning (FL) offers a privacy-preserving paradigm for this collaborative training. However, the core challenge of statistical heterogeneity across client data poses a significant challenge to the application of FL to LLMs (Zhang et al., 2024; Ye et al., 2024; Long et al., 2024; Wang et al., 2024b). A straightforward application of standard algorithms like FedAvg often leads to significant performance degradation, where the aggregated global model can underperform even a model trained solely on local data. This is due to destructive interference from conflicting client objectives, causing catastrophic forgetting and undermining both personalization and the model’s general capabilities. This highlights an urgent need for an FL framework that can effectively leverage, rather than suffer from, data heterogeneity.
Instead of forcing a single, dense model to handle conflicting data objectives, an alternative paradigm is to employ model architectures that can inherently leverage such diversity. The Mixture of Experts (MoE) (Jacobs et al., 1991) architecture presents a well-suited architectural solution to this problem. By design, MoE models route inputs to specialized "expert" sub-networks via a gating mechanism, allowing different experts to capture distinct data patterns, domains, or linguistic styles (Fedus et al., 2022; Mixtral.AI, 2024). This inherent specialization makes MoE a conceptually ideal candidate for federated settings, where experts could theoretically learn to handle the diverse data distributions from different clients. Crucially, this direction has become particularly practical and promising with the recent emergence of powerful, open-source MoE-based LLMs (DeepSeek-AI, 2025; Kimi et al., 2025; Qwen et al., 2025). The success of these models comes from an architectural tradeoff: they achieve superior performance by sparsely activating experts, but at the cost of a massive total parameter count (e.g., DeepSeek-V3 activates 37B out of total 671B parameters, Kimi-K2 activates 32B out of 1000B parameters).
Conventional FL frameworks, designed for dense models, require aggregating structurally identical parameters from all clients. Since expert activation is sparse and varies across clients based on their local data, all experts must be communicated to meet this structural requirement. This process not only prevents the system from leveraging MoE’s dynamic sparsity but also risks corrupting the specialized knowledge within pretrained experts through naive averaging. These key obstacles, particularly the risk of catastrophic forgetting that undermines the critical pretrained foundation of large-scale LLMs, pose a significant challenge to the practical application of large MoE models in personalized, privacy-preserving federated settings.
To address these challenges, we introduce FLEx (Federated LLMs with Personalized Experts), a novel federated learning framework specifically designed for pretrained MoE-based LLMs. Our approach stems from a key insight into the MoE architecture: the dense non-expert parameters (e.g., attention layers) that process all tokens serve as a repository for common knowledge, while the sparsely activated experts are ideal candidates for personalization. This natural division directly informs our decoupling strategy, whereby FLEx builds a shared foundation by exclusively aggregating the dense, common parameters, while simultaneously enabling personalization by keeping the pretrained experts frozen and introducing a novel expert grafting mechanism for clients to construct their own lightweight experts. Our main contributions are summarized as follows:
  • We propose FLEx, a novel federated learning framework that addresses the prohibitive communication costs and risk of knowledge corruption inherent in naively applying FL to MoE models. By strategically aggregating only the dense non-expert parameters while keeping the pretrained experts frozen, FLEx significantly diminishes communication overhead and reduces catastrophic forgetting of specialized world knowledge.
  • To tackle the core challenge of data heterogeneity, FLEx introduces a novel expert grafting mechanism for effective personalization. Our framework allows each client to construct a lightweight, personalized expert by pruning the frozen, pretrained experts based on local data. A jointly-tuned adaptive gating mechanism then learns to dynamically integrate this personalized expert with the shared knowledge from the base model.
  • We conduct extensive evaluations on diverse, non-IID instruction-tuning datasets. The results demonstrate that FLEx markedly outperforms established federated baselines in personalized performance while effectively preserving the general knowledge of the pretrained model, as validated by strong performance on the MMLU benchmark.
@@ 2.1 Federated Learning on Large Language Models
Federated Learning has gained significant attention as a promising method for training LLMs, addressing challenges such as decentralized data management and privacy concerns (Zhang et al., 2024; Ye et al., 2024; Wang et al., 2024b). Recent surveys (Yao et al., 2024; Ye et al., 2024) have identified important intersections between FL and LLMs, particularly in areas such as efficient foundation model training, federated fine-tuning techniques, and the collaborative potential of FL in advancing LLM development. Notable contributions in this domain include FedIT (Zhang et al., 2024), which demonstrated the utility of FL in instruction tuning tasks for LLMs. Further innovations involve personalized LLM training methods like DualLoRA (Long et al., 2024) and FDLoRA (QI et al., 2024), which combine federated learning with personalized adapter methods. Other approaches such as FRLoRA (Yan et al., ), FLoRA (Wang et al., 2024b), and FlexLoRA (Bai et al., ) introduce new aggregation strategies designed to optimize low-rank fine-tuning and allow more flexible model updates in federated settings. Additionally, selective aggregation methods, like those presented in Guo et al. (2025), further improve federated LLM training by refining the aggregation process.
Figure 1: Challenges of applying traditional federated learning to MoE-based LLMs. As shown in the right figure, in an MoE layer, a router sparsely routes tokens from different clients to a selected subset of experts. This dense aggregation scheme conflicts with MoE’s sparse activation mechanism, leading to prohibitive communication overhead and undermining the specialized knowledge of the experts.
@@ 2.2 Mixture of Experts
Mixture of Experts (Jacobs et al., 1991; Fedus et al., 2022) models activate only a subset of experts during inference or training, which significantly reduces computational costs while maintaining high performance. Recently, MoE architectures have gained attention for their scalability and efficiency in large language models (Mixtral.AI, 2024; DeepSeek-AI, 2025; Qwen et al., 2025).
However, existing federated MoE approaches (Zhan et al., 2024; Luo et al., 2025) are primarily designed for dense model architectures. Applying these methods directly to MoE-based LLMs can result in substantial computational and communication overhead. To address this limitation, we propose FLEx, a novel framework specifically designed for federated training of MoE-based LLMs.
@@ 2.3 Expert Pruning for LLMs
Model pruning is a widely used technique to improve communication and computational efficiency in deep learning. In the context of MoE-based LLMs, recent pruning methods, such as Wanda (Sun et al., 2024) and SparseGPT (Frantar & Alistarh, 2023), offer effective strategies to reduce the number of model parameters. Additionally, recent studies (Lu et al., 2024; Liu et al., 2024) have demonstrated that many experts are either unimportant or redundant during inference on specific benchmarks. However, in federated learning with MoE-based LLMs, directly pruning certain experts may result in the loss of essential foundational knowledge. To address this, we propose a method for pruning personalized side experts on local data. Furthermore, we integrate the pruned experts back into the original pretrained model to achieve an effective balance between local adaptation and global generalization.
@@ 3 Preliminary
We consider a federated learning setup with $N$ clients, each client $i\in\{1,\dots,N\}$ holding a private dataset $\mathcal{D}_{i}$. In this setup, each client independently trains a model on its local data, while the central server periodically aggregates the updates to improve the global model. A widely adopted approach, the Federated Averaging (FedAvg) algorithm, aggregates the global parameters by computing a weighted average of the local parameters:
$W^{t}=\sum_{i=1}^{N}\frac{n_{i}}{n}W_{i}^{t},$ (1)
where $n=\sum_{i=1}^{N}n_{i}$ is the total number of training samples across all clients, and $W^{t}_{i}$ and $W^{t}$ are the model weights of $i$-th client and server in $t$-th round, respectively. This approach enables the global model to leverage distributed data while maintaining privacy.
@@
The backbone of an MoE-based LLM is the Transformer architecture. A standard Transformer layer is composed of a self-attention sub-layer and a Feed-Forward Network (FFN) sub-layer. We represent the input to the $l$-th layer as a matrix $\mathbf{H}^{l-1}\in\mathbb{R}^{T\times d_{\text{model}}}$, where $T$ is the sequence length and $d_{\text{model}}$ is the hidden dimension. The layer’s operations can be expressed as:
$\mathbf{U}^{l}$ $=\text{Self-Att}\left(\text{LayerNorm}(\mathbf{H}^{l-1})\right)+\mathbf{H}^{l-1},$ (2)
$\mathbf{H}^{l}$ $=\text{FFN}\left(\text{LayerNorm}(\mathbf{U}^{l})\right)+\mathbf{U}^{l},$ (3)
In MoE models, the dense FFN sub-layer is replaced by an MoE layer, which operates on each token’s representation individually. This layer consists of $E$ expert networks (each an FFN) and a gating network, or router, that sparsely selects which experts to activate for each token. For the $j$-th token’s hidden state $u_{j}^{l}$ (a row vector from the output of attention layer $\mathbf{U}^{l}$), the MoE layer output is computed as:
$h_{j}^{l}=\left(\sum_{i=1}^{E}g_{i,j}\cdot\text{FFN}_{i}(u_{j}^{l})\right)+u_{j}^{l},$ (4)
where $g_{i,j}$ is the gate value for expert $i$ and token $j$. The gating mechanism employs a Top-K routing strategy. A router network first computes a relevance score $s_{i,j}$ for each expert $i$:
$s_{\cdot,j}=\text{Softmax}(\text{router}(u_{j}^{l})),$ (5)
The gate values are then determined by selecting the $K$ experts with the highest scores:
$g_{i,j}=\begin{cases}s_{i,j},&\text{if }i\in\text{TopK}(\{s_{1,j},\dots,s_{E,j}\},K)\\ 0,&\text{otherwise}\end{cases},$ (6)
This ensures that only a small fraction of the model’s parameters are used for each token, significantly reducing the computational overhead.
Although sparse activation provides significant computational efficiency, it presents a fundamental challenge in the context of federated learning. Standard aggregation methods like FedAvg require the communication of all model parameters. In MoE models, the total number of parameters across all experts is massive, even though only a few are active for any given input. Therefore, directly aggregating all expert layers would result in prohibitive communication costs, making the training process impractical. Furthermore, naively averaging all expert parameters risks corrupting their specialized, pretrained knowledge through destructive interference from heterogeneous client data, potentially leading to catastrophic forgetting.
@@ 4 Unlock the Power of Experts: FLEx
To address the challenges of prohibitive communication costs and knowledge corruption, we introduce FLEx (Federated LLMs with Personalized Experts), a framework designed to decouple shared knowledge from client personalization. The core strategy is to preserve the vast, specialized knowledge within pretrained experts by keeping them frozen locally, while building a robust global model by exclusively aggregating the shared, non-expert parameters. Personalization is then achieved by constructing a client-specific expert for each experts layer, not from scratch, but through a grafting mechanism. This mechanism repurposes components from the frozen, pretrained experts based on local data. This grafted expert is then integrated into the model and jointly fine-tuned with an adaptive gating mechanism. This joint training process allows the model to learn when to leverage the broad knowledge from the frozen experts and when to activate the personalized expert for client-specific tasks. The overall framework is illustrated in Figure 2.
The following subsections detail the core components of this framework: the selective aggregation strategy (§4.1), the personalized expert grafting mechanism (§4.2), and the adaptive integration process (§4.3).
@@ 4.1 Selective Aggregation in FLEx
The FLEx framework begins by partitioning the model parameters into two sets: shared non-expert parameters (e.g., self-attention layers, normalization layers) and expert parameters. Following our strategy, only the shared non-expert parameters are aggregated at the server during the federated learning process to form an updated global model.
Crucially, all original pretrained expert layers are frozen on the client side throughout the training. This design directly tackles the challenges outlined previously: it reduces communication costs by excluding the large volume of expert parameters from aggregation, and it preserves the pretrained world knowledge against destructive interference, thereby reducing catastrophic forgetting. This strategy, however, necessitates an effective mechanism for creating personalized components on the client-side. To this end, we introduce our expert grafting process.
4.2 Personalized Expert Grafting via Pruning
To address data heterogeneity and enable personalization, FLEx introduces an efficient expert grafting mechanism. Instead of training a new expert from scratch, each client constructs a lightweight, personalized expert by leveraging the rich knowledge already present in the frozen pretrained experts. This is accomplished through a targeted pruning process guided by the client’s local data.
Specifically, for each MoE layer $l$, each client $i$ identifies the single pretrained expert that best approximates the output of the full expert ensemble on its local data $\mathcal{D}_{i}$. This selection is framed as an optimization problem that minimizes the reconstruction loss between the output of a single expert and the output of the original multi-expert layer. The index of the selected expert, denoted as $e_{i,l}^{*}$, for client $i$ at layer $l$ is found by:
$e_{i,l}^{*}=\arg\min_{e\in\{1,\dots,E\}}\sum_{\mathbf{u}_{j}^{l}\in\mathcal{D}_{i}}\left\|\text{FFN}_{e}(\mathbf{u}_{j}^{l})-\sum_{k=1}^{E}g_{k,j}\cdot\text{FFN}_{k}(\mathbf{u}_{j}^{l})\right\|_{F}^{2},$ (7)
where $\mathbf{u}_{j}^{l}$ is the input to the MoE layer for the $j$-th token, $\text{FFN}_{k}$ is the $k$-th pretrained expert, $g_{k,j}$ is the original router’s gate value, and $\|\cdot\|_{F}^{2}$ denotes the squared Frobenius norm. This greedy selection process is performed once at the beginning of training for each client at each MoE layer, efficiently yielding a set of personalized experts $\{\text{FFN}_{e_{i,l}^{*}}^{l}\}_{l=1}^{L}$ that are tailored to the client’s data distribution. This collection of grafted experts forms the basis for local adaptation. These selected experts must then be integrated with the frozen pretrained experts, which we achieve through an adaptive gating mechanism detailed next.
4.3 Adaptive Integration and Fine-tuning
For each MoE layer $l$, the client’s grafted expert is integrated back into the model alongside the frozen pretrained experts. To enable the model to dynamically choose between leveraging shared knowledge and client-specific knowledge, we introduce a new, trainable gating component $\text{router}_{e,l}$ dedicated to the personalized expert at that layer. The forward pass of the modified MoE layer (Eq. 4) is updated as:
$h_{j}^{l}=\left(\sum_{k=1}^{E}g_{k,j}\cdot\text{FFN}_{k}^{l}(u_{j}^{l})\right)+g_{e,j}^{l}\cdot\text{FFN}_{e_{i,l}^{*}}^{l}(u_{j}^{l})+u_{j}^{l},$ (8)
where the gate value for the personalized expert is computed as:
$g_{e,j}^{l}=\text{sigmoid}(\text{router}_{e,l}(u_{j}^{l})).$ (9)
The sigmoid function maps the router’s output to a (0, 1) range, acting as a soft, probabilistic gate for the personalized expert. This choice was empirically validated in our ablation studies (see Section 5.3). During local training, updates are restricted to a small subset of parameters: the shared non-expert layers, the personalized expert, and its corresponding gate. All original experts and their corresponding router remain frozen. This joint training allows the model to learn when to rely on the broad knowledge of the pretrained experts and when to activate the personalized expert to handle specific local data patterns.